OpenStudy (anonymous):

You are selling tickets for a high school basketball game. Student tickets cost $3 and general admission tickets cost $5. You sell 350 tickets and collect $1,450. How many general tickets did you sell?

OpenStudy (anonymous):

Student Tickets = x General Tiickets = 350 - x do you understand why the general tickets are 350 - x?

OpenStudy (anonymous):

Hmm because 350 is the total amount of tickets sold and we need to find x?

OpenStudy (anonymous):

Right! Now set up your equation. You know student tickets are $3 & general admission tickets are $5. You also know the total collected is $1450

OpenStudy (anonymous):

3x + 5 (350)-x = 1450 solve the equation for x

OpenStudy (anonymous):

is the answer x= 100?

OpenStudy (anonymous):

Did you get this when you distributed? 3x + 1750-5x = 1450

OpenStudy (anonymous):

oops nope i see what i did wrong for that step

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

okay i must have done something wrong after that step because my answer makes no sense

OpenStudy (anonymous):

How many student tickets? How many general tickets? Be sure to label your answers...when I taught, I was pretty picky about that

OpenStudy (anonymous):

Ohhhh okay now i see what i did wrong the 5x was negative

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Okay thank you for your help!

OpenStudy (anonymous):

You're welcome